Calculate the specific heat capacity of a new alloy if a 15.4g sample absorbs 393 J when it is heated from 0.0°C to 37.6°C


1
Expert's answer
2022-03-08T08:55:16-0500

The specific heat

"c=\frac{Q}{m\Delta T}\\
=\frac{393}{15.4*(37.6-0)}=0.679\:\rm J/(g\cdot {^\circ C})"